Check whether the given fractions are equivalent:
3 / 10, 12 / 50

Grade:12

2 Answers

Harshit Singh
3 years ago
Dear Student

Given 3 / 10 and 12 / 50
We have 3 × 50 = 150
10 × 12 = 120
3 × 50 ≠ 10 × 12
Hence, 3 / 10 and 12 / 50 are not equivalent fractions

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
3/10 and 12/50 
Here, 12/50 can be reduced as 6/25
 
Taking LCM of 10 and 25 we get, 100
3×10/100 and 12×2/100
 
= 30/100 ans 24/100
 
These are not equivalent fractions.
 
Hope it helps!
Thank you